OpenStudy (anonymous):

How do i decide which reference triangle to use for sin-1 (sin 3pi/10)?

OpenStudy (noelgreco):

Do you mean\[\sin^{-1} (\sin \frac{ \pi }{ 10 })\]?

OpenStudy (noelgreco):

oops 3pi/10

OpenStudy (anonymous):

yes I do

OpenStudy (noelgreco):

First, make sure that the angle is in the domain of the inverse sine function [-2pi,2pi]. Then remember that the sin generates a number, and inverse sin generates an angle - in this case the same angle you took the sin of.